The small box provided contains a cell of emf


E


and internal resistance


r


. Using this cell, an ammeter


(with negligible resistance), a variable (but known) resistance (


R


) and a plug key connect the circuit as


shown below.


Note


the meter polarity.


(a)


Does it matter in what order the components are connected? Explain.


(b)


In this experiment


I


is the dependent and


R


is the independent variable. Is this correct? Explain.


(c)


(d)


1


1


r


R


I




E


E


Write down the equation for


E


in terms of the current,


I


, the external resistance,


R


, and the


internal resistance of the cell,


r


.


(E2.1)


Show that equation (E2.1) can be rearranged into a form suitable for plotting a straight-line


graph:




1
Expert's answer
2021-08-31T14:22:26-0400

As here question is referring a diagram but it is not attached. And question is also very splitted way, so it is not clear, hence i am taking the approximation.

As per the question,

Emf of cell = E

Internal resistance = r

Resistance of variable resistor = R


equivalent resistance of the circuit = R+r

current in the circuit "(I)=\frac{E}{r+R}"

Potential developed across the internal resistance "E_r=rI = \frac{rE}{R+r}"

Potential developed across the resistance R,

"V= IR = \frac{ER}{r+R}"

And net emf across the battery = "V_{net}=E-\frac{Er}{R+r}= \frac{ER+Er-Er}{R+r}=\frac{RE}{R+r}"